OpenStudy (exoticinspirit):

Is my work correct? Find x when f(x)=1. (This means find the input that gives an output of 1.) f(x)=2/2+2=2/4=2/2=1

Nnesha (nnesha):

y=input x=output

Nnesha (nnesha):

what is f(x) function ?

OpenStudy (exoticinspirit):

f(x)= y

OpenStudy (exoticinspirit):

the function is 2/x+2

Nnesha (nnesha):

is it \[\frac{ 2 }{ x }+2 ~~\rm or \frac{ 2 }{ x+2 }\]

OpenStudy (exoticinspirit):

the one after or (the second one)

Nnesha (nnesha):

okay cool so \[\huge\rm f(x) = \frac{ 2 }{ x+2 }\] f(x) =1 so substitute f(x) for 1 \[\rm 1=\frac{ 2 }{ x+2 }\] solve for x

OpenStudy (exoticinspirit):

!!OH Duh im finding x so i did it wrong right?

OpenStudy (exoticinspirit):

i said that x was 2 according to my calculations

Nnesha (nnesha):

yes right i don't know why u replaced `x` with `2`

OpenStudy (exoticinspirit):

x is 1

Nnesha (nnesha):

okay substitute x for 1 to check ur work \[\huge\rm 1=\frac{ 2 }{ \color{ReD}{1}+2 }\] both sides are equal ?

OpenStudy (exoticinspirit):

NO

OpenStudy (exoticinspirit):

OHHHH OK SO THEY HAVE TO EQUAL THE SAME!!! THANK YOU SO MUCH IM NEW TO THIS

OpenStudy (exoticinspirit):

:di got it let me try another #

Nnesha (nnesha):

so no x isn''t one yes right you should check ur answer if you get equal sides then yes one is right if both sides aren't equal then no

OpenStudy (exoticinspirit):

ok but there is one question where it says x=4 but when i plugged in 4 it wasnt equal

OpenStudy (exoticinspirit):



OpenStudy (exoticinspirit):

f(8) is not equal to 3 or 2/6

Nnesha (nnesha):

 bot sides are equal ?